OpenStudy (lgbasallote):

Optimization question The sum of two positive numbers is 16. What is the smallest possible value of the sum of their squares?

OpenStudy (anonymous):

uhhh 128?

OpenStudy (anonymous):

#'s are x and (16-x). Equation: y = x^2 + (16 - x)^2. Start by taking first derivative.

OpenStudy (lgbasallote):

first derivative of each?

OpenStudy (anonymous):

let's say a+a=16 now \[a^2+a^2=?\] 2a^2= y \[2a^2=y\] now \[\frac{ dy }{ dx }=2*2a=4a \] and \[\frac{ d ^{2}y }{ dx^2 }=4\] so minimum value is attained at a=0 therefore sum of their squares =0

OpenStudy (lgbasallote):

woah. one by one @jasonxx

OpenStudy (anonymous):

:)

OpenStudy (lgbasallote):

flashing whole solutions just confuse people...better to discuss the steps one by one... anyway...derivative of what @tcarroll010 ?

OpenStudy (anonymous):

the 2 numbers need not be the same, so continuing, if y = x^2 + (16 - x)^2, then y' = 2x - 2(16 - x) = 4x - 32 = 4(x-8)

OpenStudy (lgbasallote):

ironic how i just said flashing the whole solutions confuse people.....

OpenStudy (anonymous):

At x = 8 the first derivative is = 0 and it goes from negative to positive, so at x = 8, we have a minimum. Could also have taken the second derivative and see that at this point, the curve is "concave up" and hence a minimum.

OpenStudy (lgbasallote):

hmm so y' is 4(x-8)...i suppose this is equated to 0?

OpenStudy (lgbasallote):

so x = 8

OpenStudy (anonymous):

Yes. At 8 it's 0. The first derivative

OpenStudy (lgbasallote):

so if x = 8..then y = 8 too... then the sum of squares will be 128?

OpenStudy (anonymous):

Yes.

OpenStudy (lgbasallote):

thanks

OpenStudy (anonymous):

You're quite welcome!